OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through points S(–1, –4) and T(3, 4)

OpenStudy (pinksapphire):

y=2x-2

OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(6, 1) with slope 9 ^^^^^^^^^^^^^^^^^^^^^^^^ can you help me with this one?

OpenStudy (pinksapphire):

The equation for point-slope form is y-y1=m(x-x1)

OpenStudy (anonymous):

thank you so much !!!!

OpenStudy (pinksapphire):

y-1=9(x-6)

OpenStudy (pinksapphire):

Do you need any more help?

OpenStudy (anonymous):

No thats it thank you :)

OpenStudy (pinksapphire):

Okay :)